Prove that 3 + 2√5 + is irrational 
Prove that 3 + 2√5 + is irrational

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume 3 + 2√5is rational.
Then we can find co-prime x and y (y≠ 0) such that 3 + 2√5= x/y
Rearranging, we get,
2√5= x/y -3
√5= 1/2( x/y -3)
Since, x and y are integers, thus,
1/2( x/y -3)is rational number
Therefore,√5is also a rational number. But this contradicts the fact that√5is irrational.
So, we conclude that 3 + 2√5is irrational.

​Thanks